FIND Two consecutive integers such that half the smaller plus three times the larger is -32
Answer by checkley77(12569)

Let x & x+1 be the 2 integers.
x/2+3(x+1)=-32
x/2+3x+3=-32
x/2+3x=-32-3
x/2+3x=-35
3.5x=-35
x=-35/3.5
x=-10 ans for the smaller integer.
-10+1=9 ans. for the larger integer.
Proof:
-10/2+3*9=-32
-5-27=-32
-32=-32